I tried all solutions and numbers for b and c and nothing seemed to be right.

Required A Req B and C Reconstruct the income statement using the contribution margin format. JOEL COMPANY Income Statement For the Year Ended December 31, 2018 Sales revenue $ 250,000 Less: Variable costs Cost of goods sold Sales commissions Shipping and handling expenses (130,000) 25,000 (2,000) $ 143,000 Less: Fixed costs Administrative salarie Advertising expense Depreciation expense (30,000) (20,000) (24,000) Net income $ 69,000
Answer #1

Answer :

a. Reconstruct income statement using contribution margin format-

JOEL COMPANY
Income Statement
For the Year Ended 31 December, 2018
Sales revenue $ 250,000
Less : Variable Costs

Cost of goods sold

(130,000)

Sales commission

(25,000)

Shipping and handling expenses

(2,000)
Contribution Margin $ 93,000
Less : Fixed Costs

Administrative salaries

(30,000)

Advertising expense

(20,000)

Depreciation expense

(24,000)

Net Income (EBIT)

$ 19,000
b. Operating Leverage 4.89 times
c. Net income $ 28,291

Working Notes :

1. Operating Leverage = Contribution Margin ÷ EBIT

= 93,000 / 19,000 = 4.89 ( or rounded to 4.90)

2. Degree of Operating Leverage

= % Change in EBIT ÷ % Change in Sales

That is,

4.89 = % change in EBIT ÷ 10

% Change in EBIT = 4.89 × 10 = 48.9

3. % Change in EBIT (Net Income ) =

(Change in Net income ÷ Net Income ) × 100

48.9 =( Change in Net income ÷ 19,000) × 100

Change in Net income = $ 9,291** ( See below)

4. New Net Income = $ 19,000 + 9,291 = $ 28,291

** If not rounded, your in Net Income is $ 28,300.

That is, change in net income =[{(93,000/19,000)×10}×19000] ÷ 100 = 9,300